# Optimizing Product Provenance Verification using Data Valuation Methods

### v2

@@ 3.3. Data Valuation
where $\Pi(D)$ is the set of all permutations of data points in $D$, $\pi$ is a permutation sampled uniformly at random from $\Pi(D)$, and $P^{\pi}_{i}$ is the set of data points preceding instance $z_{i}$ in permutation $\pi$.
Truncated Monte Carlo Shapley Value. The exact computation of Shapley values is computationally prohibitive due to the factorial term, requiring evaluation of $2^{|S|}$ subsets. We address this shortcoming by employing the Truncated Monte Carlo Shapley estimation. This method, proposed in (Ghorbani and Zou, 2019), approximates the Shapley values by randomly sampling a limited number of subsets instead of exhaustively considering all possibilities. The key idea is that each random permutation provides an unbiased estimate of the marginal contribution of each data point. By averaging these estimates over multiple permutations, the approximation converges to the true Shapley values. The number of iterations controls the trade-off between computational cost and accuracy. Algorithm 1 shows each step in detail.
By employing the data Shapley value, particularly with the Truncated Monte Carlo approximation, we can gain valuable insights into the training data, identify influential samples, diagnose potential issues, and ultimately improve the accuracy and robustness of their GP-based timber provenance models. This data-centric approach, grounded in a strong axiomatic foundation, complements model-centric evaluations and facilitates a more efficient use of resources for data collection and model refinement.
Algorithm 1 Truncated Monte Carlo Shapley (Ghorbani and Zou, 2019).
@@
     for all $j=1$ to $N$ do
[…]
              $v_{j}^{t}=v_{j-1}^{t}$triangleright Reuse Previous Subset Performance
@@
              $v_{j}^{t}=v\left(h,P^{\pi^{t}}_{j}\cup\pi^{t}[j],T\right)$triangleright Subset Performance
@@
         $\phi_{\pi^{t}[j]}\leftarrow\frac{t-1}{t}\phi_{\pi^{t}[j]}+\frac{1}{t}\left(v_{j}^{t}-v_{j-1}^{t}\right)$triangleright Update Value
@@
3.4. Data Selection with Shapley Values
Acquiring high-quality isotope values for model training is a significant expense. Furthermore, standard data valuation techniques reveal a critical issue: not all data points contribute equally to predictive performance. In fact, some points may even be detrimental. Therefore, our objective is to strategically select a subset $D^{\prime}\subseteq D$ from the original training data $D$ that maximizes model accuracy for both forward and backward prediction tasks. We propose to leverage Shapley values computed once on the full dataset to identify data point importance. We hypothesize that by using these initial global valuations, we can efficiently identify and sequentially remove less valuable samples. Our data selection methodology consists of three distinct steps. Initially, we compute data values for all points in $D$ using the entire dataset. Subsequently, we sort the data points and remove the least valuable one according to these pre-calculated values. Finally, we evaluate the model’s performance using the test dataset. This sequential removal process continues as long as model performance improves, relying on the initial single valuation. Please refer to Algorithm 2 for a detailed algorithmic description. This simplified data selection method offers a computationally efficient approach to valuation, aiming to maximize model performance while minimizing the need for costly and potentially redundant data acquisition through a single valuation and sequential removal strategy.
@@ 3.4. Data Selection with Shapley Values
selected training dataset $D^{\prime}$ Initialize $\text{Initialize }D^{\prime}\leftarrow D$
$\phi_{i}\leftarrow\text{Compute Data Values for }D^{\prime}$triangleright Algorithm 1
@@
[…]
     $D^{\prime}\leftarrow D^{\prime}\setminus\{z_{\text{min\_i}}\}$triangleright Remove $z_{i}$ with Lowest Value
@@
4. Experiments
Our experiments are aimed at answering the below questions:
  1. (1)
  2. (2)
  3. (3)
    RQ3: Can the proposed data valuation framework enhance outcomes in cases involving data imputation for missing or noisy data? (Section 4.5)
  4. (4)
    RQ4: Can data selection methods based on data valuation improve the performance of both directions in SIRA? (Section 4.6)
  5. (5)
    RQ5: Given a specific model and data valuation framework, what level of granularity is optimal for effective data selection? (Section 4.7)
  6. (6)
    RQ6: Do different genera and species within the dataset exhibit varying data Shapley values? Can we identify the most and least important genera or species within the dataset? (Section 4.8)
4.1. Datasets
We utilized data from two datasets of the genus Quercus, collected from various regions worldwide. Two broad datasets were combined in this study with the first dataset comprising tree samples distributed globally, while the second dataset was focuses specifically on European countries. Stable isotope ratio measurements were performed following the protocols outlined in (Watkinson et al., 2020; Boner et al., 2007).
Figure 1. Distribution of data Shapley values (top: backward model, bottom: forward model) (Section 4.3)
(a) Error plot for Gaussian process regression
(b) Error plot for random forest
Figure 2. Effect of model architecture on data valuation framework (Section 4.4)
(a) Error plot for median-based imputation
(b) Error plot for listwise deletion
(c) Data Shapley value distribution for median-based imputation
(d) Data Shapley value distribution for listwise deletion
Figure 3. Enhancing missing data strategies through data valuation (Section 4.5)
(a) Error plot for removing low-value data points
(b) Error plot for removing high-value data points
(c) Data Shapley value distribution per species (Section 4.8)
Figure 4. Optimal Granularity in data selection (Section 4.7) and Data Shapley value distribution per species (Section 4.8)
(a) Error plot for backward direction
(b) Error plot for random forest on forward direction
(d) Data Shapley rank overlap for forward vs backward direction
Figure 5. Random Forest-based data valuation on USA only data (Section 4.6)
(a) Error plot for backward direction
(b) Error plot for random forest on forward direction
(d) Data Shapley rank overlap for forward vs backward direction
Figure 6. Random Forest-based data valuation on Europe only data (Section 4.6)
4.2. Data Valuation and Selection
As described in Section 3, the objective of data valuation is to strategically select subsets of the dataset that maximize model performance. Instead of solely selecting the most valuable data points, we implement an iterative approach. Given a dataset $D$ consisting of data points $(z_{1},z_{2},\dots,z_{i})$ and their corresponding data Shapley values $(\phi_{1},\phi_{2},\dots,\phi_{i})$, we conduct experiments by removing the low value data points $z_{k}$, where the data Shapley value $\phi_{k}$ satisfies $\phi_{k}<\phi_{j}\quad\forall j\neq k$, or equivalently, $\phi_{k}=\min_{j}\phi_{j}$. Furthermore, we extend our experiments by investigating the opposite case: removing the high-value data points $z_{m}$, where the Shapley value $\phi_{m}$ satisfies $\phi_{m}>\phi_{j}\quad\forall j\neq m$, or equivalently, $\phi_{m}=\max_{j}\phi_{j}$. We then evaluate model performance and analyze how the data valuation framework influences the overall model behavior. We report the root mean square error (RMSE) for all performance plots, defined as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$, where $y_{i}$ and $\hat{y}_{i}$ represent the actual and predicted multivariate outputs involving variables $\mathcal{X}$ or $\mathcal{Y}$, on both forward and backward directions.
4.3. RQ1: Role of Data Shapley Values in SIRA
4.4. RQ2: Influence of Model Architecture
We conducted two types of experiments, removing both high value and low value data points, as described in Section 4.2. We present the error plots showing the effect of removing these two types of data points for Gaussian process regression and random forest models. In both models, the RMSE increases (indicating performance degradation) when data points with high Shapley values are removed. Similarly, the RMSE decreases (indicating performance improvement) or remains stable when data points with low Shapley values are removed (Figure 2(a) and Figure 2(b)). Figure 2(c) illustrates one such case, where performance improves by removing specific data points (marked as red x on the map).
We further investigate whether the Shapley values from two different models agree with each other. We present three types of rank comparison plots in Figure 2(d-f). All three representations demonstrate a high degree of agreement between the two models in this experiment, indicating that data value ranks can remain consistent across different model architectures.
4.5. RQ3: Data Valuation to support Missing Data Imputation
We conduct experiments with two types of approaches for handling missing data: (i) median imputation, where missing values are replaced with the median of the dataset, and (ii) listwise deletion, where data points containing missing values are excluded entirely from the training process. Additionally, we perform data selection experiments by removing both high-value and low-value data points and comparing the corresponding performance rankings.
We conducted experiments involving the removal of both high-value and low-value data points, as described in Section 4.2. The error plots (Figures 3(a) and 3(b)) illustrate the effects of removing these data points under both missing data handling strategies. For both strategies, the RMSE increases (indicating performance degradation) when data points with high Shapley values are removed. Conversely, the RMSE decreases (indicating performance improvement) or remains stable when data points with low Shapley values are removed.
We observe significant performance improvements when low-value data points are removed from both strategies, with the improvement being more pronounced for median imputation (26.66%) compared to listwise deletion (5.88%). This result demonstrates the effectiveness of data valuation as a strategy for improving missing data handling. Moreover, the Shapley values for the median imputation strategy exhibit higher magnitudes compared to listwise deletion (Figures 3(c) and 3(d)), which aligns with the greater performance improvements observed for median imputation.
4.6. RQ4: Data Selection Methods for Forward and Backward directions in SIRA
USA-only data: We present the error plots showing the effect of removing high-value versus low-value data points on the performance of the random forest model. In both directions, the RMSE increases (indicating performance degradation) when data points with high Shapley values, are removed. Similarly, the RMSE either decreases (indicating performance improvement) or remains stable when data points with low Shapley values, are removed (Figures 5(a) and 5(b)). Figure 5(c) illustrates a specific case where performance improves following the removal of certain data points, marked as red x on the map.
We further investigate whether the Shapley values from both directions show agreement. To this end, we present three types of rank comparison plots in Figure 5(d-f). All three representations indicate a high degree of agreement between the two directions, suggesting that data value ranks remain relatively consistent across different directions of SIRA.
Europe-only data: We conduct the same set of experiments and analyses on the Europe-only dataset. Similar patterns of performance improvement and deterioration are observed when low-value and high-value data points are removed, respectively (Figure 6). In the backward direction, performance improved by 5.08%, while in the forward direction, it improved by 7.81%.
We also compared the Shapley value ranks of the data points. For the Europe-only dataset, the similarities were less ideal than other cases, indicating that the ranking of data importance can differ across directions in different datasets. Figure 6(d-f) illustrates the corresponding plots for this case.
4.7. RQ5: Optimal Granularity in Data Selection
Here, we analyze the appropriate level of granularity for data selection to maximize performance when using a specific model and a data valuation framework. Instead of removing one data point at a time, we consider clusters of fixed distances (in kilometers) and remove all data points within that distance if a data point is selected for removal during the data selection process. We performed this location-based data selection for both high value (Figure 4(b)) and low value (Figure 4(a)) data points to understand the impact on model performance.
Consistent with the results of previous experiments, we observe performance improvements when low-value data points are removed. Furthermore, the cluster-based removal approach enhances the model’s performance more effectively than the one-by-one removal approach. This finding suggests that certain regions within the dataset are not essential for the best-performing model, or the improvements may be attributed to errors in the data extraction pipelines affecting specific regions.
4.8. RQ6: Species-Specific Data Shapley Values and Their Implications
For this experiment, we explore the variations in data Shapley values across different genera and species, aiming to identify the most and least significant contributors within the dataset. We present a sample distribution plot of data Shapley values for individual species within one dataset (Figure 4(c)). The plot demonstrates that certain species exhibit significantly higher data Shapley values compared to others, indicating their greater contribution to the model’s performance.
5. Description of Deployment
As described earlier, our approach is part of a deployed ML pipeline that supports provenance identification for timber products. Our framework is in use by European enforcement agencies to assist claims verification of wood suspected of being in violation of sanctions against Russia by demonstrating that a claimed harvest location other than Russia is not viable. See coverage of our work from the New York Times (Nazaryan, 2024). Due to confidentiality reasons, we have presented results on a Quercus dataset.
6. Conclusion and Future Work
Data valuation based on Shapley values demonstrates promising results in SIRA analytics. Additionally, region-based data selection has been shown to further enhance model performance by effectively removing low-value data clusters. Our analysis highlights that the greatest performance improvements are observed when low-value data points with higher absolute Shapley values are removed, and similarly, larger performance drops occur when high-value data points with higher absolute Shapley values are excluded. This demonstrates that data values inferred by our approach are meaningful and can be contextualized for use in product provenance verification. Future work is aimed at generalizing data valuation methodologies to span more parts of natural resource supply chains.

### v3

@@ 3.2 Data Valuation
where $\Pi(D)$ is the set of all permutations of data points in $D$, $\pi$ is a permutation sampled uniformly at random from $\Pi(D)$, and $P^{\pi}_{i}$ is the set of data points preceding instance $z_{i}$ in permutation $\pi$. This permutation-based form is equivalent to the subset definition and underlies the Monte Carlo approximation (Ghorbani and Zou 2019; Kwon and Zou 2021). In practice, we use the permutation form primarily for intuition, while our actual computations employ the subset-sampling TMC-Shapley method.
Truncated Monte Carlo Shapley Value. Because the exact computation of Shapley values is computationally prohibitive, (Ghorbani and Zou 2019) propose their approximation by randomly sampling a limited number of subsets instead of exhaustively considering all possibilities (see Algorithm 1 in the extended version). The key idea is that each random permutation provides an unbiased estimate of the marginal contribution of each data point, convergence achievable in practice in $O(n)$ permutations (typically around $3n$) for an $n$-point dataset. Generally, Monte Carlo estimators exhibit variance that decreases proportional to $1/\sqrt{m}$ as the number of samples $m$ increases. However, in SIRA settings, each marginal evaluation entails retraining a Gaussian process with $\mathcal{O}(N^{3})$ time complexity, creating a pronounced trade-off between computational feasibility and valuation precision. This tension motivates careful calibration of the permutation budget to balance estimator fidelity against real-world processing constraints.
Iterative Data Selection with Shapley Values. To strategically select a subset $D^{\prime}\subseteq D$ from the original training data $D$ that maximizes model accuracy for both forward and backward prediction tasks, we propose to leverage Shapley values computed once on the full dataset to identify data point importance. We hypothesize that by using these initial global valuations, we can efficiently identify and sequentially remove less valuable samples. Our data selection methodology consists of three distinct steps. Initially, we compute data values for all points in $D$ using the entire dataset. Subsequently, we sort the data points and remove the least valuable one according to these pre-calculated values. Finally, we evaluate the model’s performance using the test dataset. This sequential removal process continues as long as model performance improves, relying on the initial single valuation (see Algorithm 2 in the extended version).
Beta Shapley. Building upon the foundational principles of the Shapley value, Beta Shapley offers a flexible generalization for data valuation by recognizing that the standard Shapley value’s uniform weighting of a data point’s marginal contribution across all subset sizes is not always optimal. In many data valuation tasks, the primary objective is to rank data points rather than precisely distribute the total model performance gain, which makes the strict efficiency axiom less critical. Beta Shapley therefore relaxes this axiom and introduces a weighted-average framework in which weights are governed by a Beta distribution. This enables a more fine-grained valuation by assigning different levels of importance to a data point’s marginal contribution depending on the cardinality of the subset it is added to. For instance, by selecting appropriate parameters for the Beta distribution, one can prioritize the contributions made to smaller subsets (see Algorithm 3 in the extended version). We include Beta Shapley primarily to provide conceptual background within the broader family of Shapley-based data valuation methods. Our focus remains on TMC-Shapley, given its superior scalability and lower computational variance in practice.
4 Experiments
We utilized data from two datasets of the genus Quercus, collected from various regions worldwide. Two broad datasets were combined in this study with the first dataset comprising tree samples distributed globally ($N=491$), while the second dataset was focuses specifically on European countries ($N=287$). Stable isotope ratio measurements were performed following the protocols outlined in (Watkinson et al. 2020; Boner et al. 2007). Our experiments are aimed at answering the below questions:
  1. 1.
  2. 2.
  3. 3.
    RQ3: How does Shapley-based data selection compare against naive or exhaustive baselines? (Section 4.3)
  4. 4.
    RQ4: Can the proposed data valuation framework enhance outcomes in cases involving data imputation for missing or noisy data? (Section 4.4)
  5. 5.
    RQ5: Can data selection methods based on data valuation improve the performance of both directions in SIRA? (Section 4.5)
  6. 6.
    RQ6: Given a specific model and data valuation framework, what level of granularity is optimal for effective data selection? (Section 4.6)
  7. 7.
    RQ7: Do different genera and species within the dataset exhibit varying data Shapley values? Can we identify the most and least important genera or species within the dataset? (Section 4.7)
Figure 1: Distribution of data Shapley values, green and red lines represent the mean and the zero, respectively (top: backward model, bottom: forward model); see Section 4.1.
(a) Error plot for Gaussian process regression
(b) Error plot for random forest
Figure 2: Effect of model architecture on data valuation framework ($N=87$); see Section 4.2.
(a) Error plot for median-based imputation
(b) Error plot for listwise deletion
(c) Data Shapley value distribution(median-based imputation)
(d) Data Shapley value distribution for listwise deletion
Figure 3: Enhancing missing data strategies through data valuation ($N=491$); see Section $4.$
(a) Error plot for backward direction
(b) Error plot for random forest on forward direction
(d) Data Shapley rank overlap for forward vs backward direction
Figure 4: Random Forest-based data valuation on USA only data ($N=87$); see Section 4.5.
(a) Error plot for backward direction
(b) Error plot for random forest on forward direction
(d) Data Shapley rank overlap for forward vs backward direction
Figure 5: Random Forest-based data valuation on Europe only data ($N=287$); see Section 4.5.
(a) Error plot for removing low-value data points
(b) Error plot for removing high-value data points
(c) Data Shapley value distribution per species (Section 4.7)
Figure 6: Optimal Granularity in data selection ($N=491$); see Section 4.6 and Data Shapley value distribution per species (see Section 4.7).
4.1 RQ1: Role of Data Shapley Values in SIRA
4.2 RQ2: Influence of Model Architecture
We conducted two types of experiments, removing both high value and low value data points. We present the error plots showing the effect of removing these two types of data points for Gaussian process regression and random forest models. In both architectures, the RMSE increases (indicating performance degradation) when data points with high Shapley values are removed. Conversely, the RMSE decreases (indicating performance improvement) or remains stable when data points with low Shapley values are removed (Figure 2(a) and Figure 2(b)). Figure 2(c) illustrates one such case, where performance improves by removing specific data points (marked as red x on the map). Beyond accuracy, we should point out that random forests, while lacking the theoretically principled uncertainty quantifications of GPR, are much more scalable. We further investigate whether the Shapley values from the two architectures agree with each other. Figures 2(d-f) present three types of rank comparison plots, all of which demonstrate a high degree of agreement between the two models. This result suggests that data value ranks remain consistent across architectures, reinforcing the robustness of our valuation framework even when computational constraints necessitate different modeling choices.
4.3 RQ3: How does Shapley-based data selection compare against naive or exhaustive baselines?
To contextualize the effectiveness of our Shapley-based data valuation framework, we compare it against two natural baselines—random removal of data points and leave-one-out (LOO) removal. Using USA-only data (which already captures the main performance trends observed across our study), Table 1 shows that our Truncated Monte Carlo (TMC) Shapley approach delivers the largest improvement in predictive accuracy, reducing RMSE by 0.8459 compared to the initial model, nearly four times greater than the improvement achieved with random removal and significantly outperforming LOO.
Method Initial RMSE Best RMSE Points Removed $\Delta$ RMSE
Random Removal 3.6101 3.4208 2 0.1893
Remove Low-Value with LOO 3.6101 3.2435 24 0.3665
Remove Low-Value with TMC-Shapley (ours) 3.6101 2.8076 22 0.8025
Table 1: Comparison of data selection strategies. Shapley-based removal consistently outperforms random removal and achieves greater RMSE reduction than LOO, while being more theoretically grounded and computationally efficient. (Section 4.3)
4.4 RQ4: Data Valuation to support Missing Data Imputation
A critical but often underappreciated issue is that imputation strategies implicitly encode assumptions about isotopic stationarity, e.g., median imputation presumes that the conditional distribution of missing values is homogeneous across all spatial and taxonomic contexts (an assumption that may be violated under latitudinal gradients or localized environmental variation). Similarly, listwise deletion assumes that missingness is independent of the isotopic signal (which is rarely true in practice, as remote sites are both more difficult to sample and may exhibit distinct isotopic patterns). These hidden assumptions complicate the interpretation of downstream model performance and underscore the need for principled data valuation frameworks to mitigate imputation-induced biases. To explore this, we conduct experiments with two common approaches for handling missing data: (i) median imputation, replacing missing values with the dataset median, and (ii) listwise deletion, excluding data points containing missing values from training; in addition, we perform data selection experiments by removing both high-value and low-value data points and examining the resulting performance rankings.
We conducted experiments involving the removal of both high-value and low-value data points. The error plots (Figures 3(a) and 3(b)) illustrate the effects of removing these data points under both missing data handling strategies. For both strategies, the RMSE increases (indicating performance degradation) when data points with high Shapley values are removed. Conversely, the RMSE decreases (indicating performance improvement) or remains stable when data points with low Shapley values are removed. We observe significant performance improvements when low-value data points are removed from both strategies, with the improvement being more pronounced for median imputation (26.66%) compared to listwise deletion (5.88%). This result demonstrates the effectiveness of data valuation as a strategy for improving missing data handling. Moreover, the Shapley values for the median imputation strategy exhibit higher magnitudes compared to listwise deletion (Figures 3(c) and 3(d)), which aligns with the greater performance improvements observed for median imputation.
4.5 RQ5: Data Selection Methods for Forward and Backward directions in SIRA
USA-only data: We present the error plots showing the effect of removing high-value versus low-value data points on the performance of the random forest model. In both directions, the RMSE increases (indicating performance degradation) when data points with high Shapley values, are removed. Similarly, the RMSE either decreases (indicating performance improvement) or remains stable when data points with low Shapley values, are removed (Figures 4(a) and 4(b)). Figure 4(c) illustrates a specific case where performance improves following the removal of certain data points, marked as red x on the map.
We further investigate whether the Shapley values from both directions show agreement. To this end, we present three types of rank comparison plots in Figure 4(d-f). All three representations indicate a high degree of agreement between the two directions, suggesting that data value ranks remain relatively consistent across different directions of SIRA. Similar results for the Europe-only dataset are shown in Fig. 5(a-f).
4.6 RQ6: Optimal Granularity in Data Selection
Here, we analyze the appropriate level of granularity for data selection to maximize performance when using a specific model and a data valuation framework. Instead of removing one data point at a time, we consider clusters of fixed distances (in kilometers) and remove all data points within that distance if a data point is selected for removal during the data selection process. We performed this location-based data selection for both high value (Figure 6(b)) and low value (Figure 6(a)) data points to understand the impact on model performance. Consistent with the results of previous experiments, we observe performance improvements when low-value data points are removed. Furthermore, the cluster-based removal approach enhances the model’s performance more effectively than the one-by-one removal approach.
4.7 RQ7: Species-Specific Data Shapley Values and Their Implications
For this experiment, we explore the variations in data Shapley values across different genera and species, aiming to identify the most and least significant contributors within the dataset. We present a sample distribution plot of data Shapley values for individual species within one dataset (Figure 6(c)). The plot demonstrates that certain species exhibit significantly higher data Shapley values compared to others, indicating greater contribution to the model’s performance.
5 Deployment Details
WFID has operationalized the SIRA approach in this paper to trace the geographic origin of forest products by guiding the collection of high-risk reference samples (Norman 2023), analyzing them in certified laboratories (Fera Science 2025), and deploying trained geographic origin models via the World Forest ID Evaluation Platform (World Forest ID 2024b), which is used by regulators, certification bodies, auditors, and companies to verify sourcing claims across timber and other EU Deforestation Regulation-relevant commodities such as soy, cacao (World Forest ID 2024a), and coffee. In a recent timber market study (Greenfield 2025; World Forest ID 2025), corporate partners used the WFID platform to assess sourcing claims for 59 wood products, focusing on birch due to concerns over sourcing from sanctioned regions, where SIRA and spatial models were used to validate or refute claimed harvest locations. WFID has also supported enforcement efforts, helping identify over 260 tons of allegedly illegal timber (Speed 2024) in Belgium and supporting at least nine additional ongoing investigations.
6 Conclusion and Future Work
Our data valuation methods demonstrate promising results in SIRA analytics, and region-based data selection further enhances performance by removing low-value data clusters. The greatest gains are observed when low-value data points with higher absolute Shapley values are removed, while the largest drops occur when high-value points are excluded, showing that the inferred values are meaningful for product provenance verification. Our analyses further indicate that Shapley-based valuation captures broader spatial and distributional informativeness across regions and species, rather than merely filtering noisy samples, reinforcing its robustness and interpretability for provenance modeling. Future work will generalize these methods across natural resource supply chains, incorporating cross-modal data and exploring greater model scalability and robustness.
Acknowledgments
This paper is based upon work supported by the NSF under Grant No. CMMI-2240402. JT is supported by the Swedish Foundation for Strategic Environmental Research MISTRA (Utmana program). Ruoxi Jia and the ReDS lab also acknowledge support from the National Science Foundation through grants IIS-2312794, IIS-2313130, and OAC-2239622. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
Appendix A Technical Appendix
Algorithm 1 Truncated Monte Carlo Shapley (Ghorbani and Zou 2019).
@@ Appendix A Technical Appendix
  for all $j=1$ to NN do
   if  $\text{ if }\left|v(D)-v_{j-1}^{t}\right|<\text{ Performance Tolerance}$ then
     $v_{j}^{t}=v_{j-1}^{t}$ $\triangleright$ Reuse Previous Subset Performance
@@
     $v_{j}^{t}=v\left(h,P^{\pi^{t}}_{j}\cup\pi^{t}[j],T\right)$ $\triangleright$ Subset Performance
@@
   $\phi_{\pi^{t}[j]}\leftarrow\frac{t-1}{t}\phi_{\pi^{t}[j]}+\frac{1}{t}\left(v_{j}^{t}-v_{j-1}^{t}\right)$ $\triangleright$ Update Value
@@
selected training dataset $D^{\prime}$ Initialize $D\text${Initialize $^{\prime}\leftarrow D$
$\phi_{i}\leftarrow\text{Compute Data Values for }D^{\prime}$ $\triangleright$ Algorithm 1
@@
  $\text{min\_i}\leftarrow\arg\min_{i}\phi_{i}$ $\triangleright$ Find $z_{i}$ with Lowest Value
  $D^{\prime}\leftarrow D^{\prime}\setminus\{z_{\text{min\_i}}\}$ $\triangleright$ Remove $z_{i}$ with Lowest Value
@@
Algorithm 3 Efficient Monte Carlo Beta Shapley.
1:Datasets $D=\{z_{1},\dots,z_{n}\}$, $T$, function $h$, performance metrics $v$, Beta parameters $\alpha,\beta$.
2:Beta Shapley values $\psi_{i},\forall i\in[n]$.
3:Initialize values $\psi_{i}=0,\forall i\in[n]$.
4:Initialize iteration counter $B=1$.
5:for all $k=1$ to $n$ do $\triangleright$ Pre-compute normalized weights
6:  $\tilde{w}_{k}\leftarrow n\binom{n-1}{k-1}^{-1}\frac{\text{Beta}(k+\beta-1,n-k+\alpha)}{\text{Beta}(\alpha,\beta)}$
7:end for
8:while not converged do
9:  for all $j=1$ to $n$ do
10:   Sample a cardinality $k$ uniformly from $\{1,\dots,n\}$.
11:   Sample a subset $S\subset D\setminus\{z_{j}\}$ with |S|=k1|S|=k-1 uniformly at random.
12:   $\Delta\leftarrow v(h,S\cup\{z_{j}\},T)-v(h,S,T)$. $\triangleright$ Calculate marginal contribution
13:   $\psi_{j}\leftarrow\frac{B-1}{B}\psi_{j}+\frac{1}{B}\cdot\tilde{w}_{k}\cdot\Delta$. $\triangleright$ Update value with Beta weight
14:  end for
15:  $B\leftarrow B+1$.
16:end while
17:return $\psi_{i},\forall i\in[n]$